A child compresses his spring gun 1 cm. If 3 mJ of energy are stored in the spring, what is the spring

constant?


1
Expert's answer
2021-02-07T19:11:06-0500

By the definition of the potential elastic energy, we have:


"PE=\dfrac{1}{2}kx^2,""k=\dfrac{2PE}{x^2}=\dfrac{2\cdot3\cdot10^{-3}\ J}{(0.01\ m)^2}=60\ \dfrac{N}{m}."